OpenStudy (anonymous):

the measures, in degrees, of the interior angles of a 15-sided polygon are consecutive integers. The number of degrees in the smallest angle is?

OpenStudy (dumbcow):

sum=180(n-2) sum=180(13) sum=2340 then the angles make a sequence n+(n+1) + ...(n+14) = 2340 15n +(1+2+..14) = 2340 15n + (14/2)(1+14) = 2340 15n + 105 = 2340 n = 149